OpenStudy (anonymous):

what is X if -4<3x+2

OpenStudy (anonymous):

\[-4 <3x+2 \le 14\]substract 2\[-6 < 3x \le 12\]divide by 3\[-2 < x \le 4\]

OpenStudy (anonymous):

Thank you so much!